OpenStudy (anonymous):

Find the sum. 6x + 19x = is it 25x or 25x^2

OpenStudy (mathstudent55):

is 6 books + 19 books = 25 books or = 25 books squared?

OpenStudy (anonymous):

lol 25

OpenStudy (mathstudent55):

x can only change to x^2 if it is _multiplied_ by x: x * x = x^2 x + x = 2x x^2 + x^2 = 2x^2

OpenStudy (mathstudent55):

x^2 + x = x^2 + x (it's already simplified)

OpenStudy (mathstudent55):

6x + 19x = 25x

OpenStudy (anonymous):

ohkay thank youuuuu .

OpenStudy (mathstudent55):

wlcm

OpenStudy (anonymous):

it is 25x;

OpenStudy (anonymous):

factorise 6x +19x = x(6+19) =x25 =25X